OpenStudy (anonymous):

Steps to convert 330 minutes to hours properly, i feel im doing it wrong.

OpenStudy (anonymous):

I got 5.5 hours.

OpenStudy (anonymous):

330 divided by 60

OpenStudy (anonymous):

So you're right!

OpenStudy (anonymous):

Oh great! Thanks! i was just checking

ganeshie8 (ganeshie8):

\(\large 330 \text{minutes}\) \(\large 330 \text{minutes} \times \frac{\text{1 hour}}{\text{60 minutes}} \) \(\large 330 \cancel{\text{minutes}} \times \frac{\text{1 hour}}{60 \cancel{\text{ minutes}}} \)

OpenStudy (anonymous):

Oh thank you for the visual! This helped me get it better.

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:D